# Susceptibilities of rotating quark matter in Fourier-Bessel basis

### v1

@@ III.1 Meson susceptibilities
$G_{\sigma}=G_{\pi},\quad G_{\eta}=G_{\delta}=g_{s}-g_{d},$ (49)
and the one-loop results are given by
$\begin{split}&\chi_{\eta}^{({\rm 1-loop})}\left(r;M(r)\right)=\chi_{\pi}^{({\rm 1-loop})}\left(r;M(r)\right),\\ &\chi_{\sigma}^{({\rm 1-loop})}\left(r;M(r)\right)=\chi_{\delta}^{({\rm 1-loop})}\left(r;M(r)\right)\\ &=N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\left[\frac{p_{l,k}^{2}+p_{z}^{2}}{E^{3}}\left(f(p_{z};\Lambda)-2n_{F}(E)\right)-\frac{2M^{2}}{TE^{2}}n_{F}^{\prime}(E)\right],\end{split}$ (50)
with $n_{F}^{\prime}(E)=dn_{F}/dE$. As with the one-loop result for the pion susceptibility, the resumed meson susceptibilities $\chi_{\alpha}$ also become independent of rotation $\Omega$ at zero temperature. The rotational effect at finite temperature will be specifically analyzed through numerical calculations in the next section.
III.2 Other susceptibility functions
@@ III.2.1 Topological susceptibility
${\cal L}_{\theta}=\theta Q,\quad Q=\frac{g^{2}}{64\pi^{2}}\epsilon^{\mu\nu\rho\sigma}G_{\mu\nu}^{a}G_{\rho\sigma}^{a},$ (51)
@@
$\int d^{4}x\chi_{\rm top}(x)=\frac{\partial^{2}\Gamma_{\rm QCD}}{\partial\theta^{2}}\Biggl{|}_{\theta=0}.$ (52)
By further evaluating the right-hand side in Eq. (52) with the $\theta$-term, the topological susceptibility is expressed by the two-point correlation function of the topological operator,
$\chi_{\rm top}=\int d^{4}x^{\prime}\langle T_{\tau}Q(x)Q(x^{\prime})\rangle.$ (53)
However, this expression may not be directly manageable within the effective model analysis, as the gluon degrees of freedom are not explicitly included. Nevertheless, by using the WT identities associated with the $U(1)_{A}$ anomaly, it is possible to relate the topological susceptibility to the quark degree of freedom, specifically through the meson susceptibility. This allows the effective model analysis to be applied in the evaluation of the topological susceptibility.
Under the $U(1)_{A}$ transformation, the $\theta$-dependence in the QCD Lagrangian is transferred into the mass term through the $U(1)_{A}$ anomaly (for the detailed procedure, see, e.g., Ref. [48, 50]):
$-m_{0}\bar{q}q+\theta Q\to-m_{0}\bar{q}\exp\left[i\gamma_{5}\theta/2\right]q.$ (54)
Following the definition of the topological susceptibility in Eq. (52) with the $\theta$-dependent mass term in Eq. (54), the topological susceptibility can be expressed as the difference between the meson susceptibilities [48, 49, 50, 45, 46]:
$\chi_{\rm top}\left(r;M(r)\right)=\frac{m_{0}^{2}}{4}\left[-\chi_{\pi}\left(r;M(r)\right)+\chi_{\eta}\left(r;M(r)\right)\right].$ (55)
Here, we have used the WT identity (30). Note that, since the $SU(2)_{A}$ and $U(1)_{A}$ symmetries are intact even in the rotational cylindrical frame, the analytical expression in Eq. (55) is identical to that in flat spacetime without rotation. However, the topological susceptibility implicitly acquires the $r$-dependence and the rotational effect through the meson susceptibilities at finite temperature. By using the meson susceptibilities obtained in the previous subsection, the topological susceptibility can be evaluated within the NJL model under rotation.
III.2.2 Baryon number susceptibility
@@
$\begin{split}\int d^{4}x\chi_{B}(x)&=\frac{\partial^{2}\Gamma}{\partial\mu^{2}}\Biggl{|}_{\mu=0}=\int d^{4}x\int d^{4}x^{\prime}\langle T_{\tau}j_{B}(x)\,j_{B}(x^{\prime})\rangle.\end{split}$ (56)
@@
$\chi_{B}\left(r;M(r)\right)=N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\left\{-2n_{F}^{\prime}(E)\right\}.$ (57)
@@ III.2.3 Moment of inertia
$\begin{split}\int d^{4}x\chi_{\Omega}(x)&=\frac{\partial^{2}\Gamma}{\partial\Omega^{2}}\Biggl{|}_{\Omega=0}=\lim_{\Omega\to 0}\left[\int d^{4}x\int d^{4}x^{\prime}\langle T_{\tau}j_{\Omega}(x)\,j_{\Omega}(x^{\prime})\rangle\right].\end{split}$ (58)
@@
$\chi_{\Omega}\left(r;M(r)\right)=N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\left\{-2n_{F}^{\prime}(E)\right\}\left(l+\frac{1}{2}\right)^{2}.$ (59)
@@
IV Numerical evaluation of susceptibilities
@@
$\frac{\delta\,\Gamma\left(M\right)}{\delta M}=0.$ (60)
@@
$\begin{split}\Gamma\left(M\right)T&=-\int d^{3}x\frac{(M-m_{0})^{2}}{2(g_{s}+g_{D})}+N_{c}N_{f}\int d^{3}x\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\\ &\quad\times\left[Ef(p_{z};\Lambda)+2T\ln\left(1+e^{-\beta\left\{E-\Omega\left(l+\frac{1}{2}\right)\right\}}\right)\right].\end{split}$ (61)
@@
$M\left(r\right)=m_{0}+(g_{s}+g_{d})N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\frac{M}{E}\left(f^{(l,k)}(p_{z};\Lambda)-2n_{F}(E)\right).$ (62)
@@
Using the dynamical quark mass in Eq. (62), one can verify the validity of the WT identity in Eq. (30) within the NJL model under rotation. From the dynamical quark mass, the quark condensate can be expressed as
$\langle\bar{q}q\rangle=-N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\frac{M}{E}\left(f(p_{z};\Lambda)-2n_{F}(E)\right).$ (63)
Substituting this into the pion susceptibility in Eq. (46), one finds that the WT identity is satisfied:
$\begin{split}\chi_{\pi}&=\frac{-\langle\bar{q}q\rangle}{M+G_{\pi}\langle\bar{q}q\rangle}=\frac{-\langle\bar{q}q\rangle}{m_{0}}.\end{split}$ (64)
When formulating the quark propagator within the local density approximation, we have analytically neglected the $O(\partial_{r}M/M^{2})$ term in Eq. (35). Therefore, our numerical analysis is valid provided that the $r$-dependent quark mass satisfies the following condition:
$\frac{|\partial_{r}M|}{M^{2}}\ll 1.$ (65)
@@
As we mentioned below Eq. (44), at $T=0$ the dynamical quark mass is unaffected by finite $\Omega$ effects, though it still exhibits $r$-dependence. To numerically examine the $r$-dependence of the dynamical quark mass in the cylindrical frame at fixed temperatures, we present panel (a) of Fig. 1. At $T=0$, the dynamical quark mass remains constant up to approximately $r=0.85\,R$ but begins to decrease sharply near the boundary to satisfy the boundary condition. Moving on to finite temperature systems, the dynamical quark mass is affected by the rotational effect even for $r<0.85R$, and it decreases with increasing $r$.
To confirm the validity of the local density approximation, we plot $|\partial_{r}M|/M^{2}$ in panel (b). This result shows that $|\partial_{r}M|/M^{2}\ll 1$ is satisfied for approximately $r<0.9\,R$ even at high temperatures across the pseudocritical temperature with the finite $\Omega$. This implies that the local density approximation is valid up to $r=0.9\,R$ even when the rotational effect is included at the high temperatures. However, in the region near the boundary for $0.9\,R<r<R$, $|\partial_{r}M|/M^{2}$ increases sharply, and the validity of the local density approximation becomes unreliable. Based on this evaluation, our analysis in the following subsections will focus on the region $r<0.9\,R$.
@@
IV.1 Meson and topological susceptibilities
In Fig. 3, we numerically evaluate the meson susceptibilities under rotation. Panel (a) illustrates its $r$-dependence at fixed temperature, both with and without the rotational effect, clearly showing the difference in the $r$-dependence between the scalar meson susceptibilities $(\chi_{\sigma,\delta})$ and pseudoscalar meson susceptibilities $(\chi_{\pi,\eta})$. This difference comes from the analytical one-loop results evaluated in Eq. (50). Looking at the details of each meson susceptibilities, the sigma meson susceptibility, which corresponds to the isosinglet scalar channel, remains approximately constant in the relevant region ($r<0.9\,R$) at zero temperature. At finite temperature, its magnitude becomes larger, but it still exhibits constant behavior. When including the rotational effect, the sigma meson susceptibility increases with $r$ and then decreases near the boundary in order to satisfy the boundary condition. Similarly, the delta meson susceptibility, which corresponds to the isotriplet scalar channel, shows a qualitatively similar behavior to that of the isosinglet channel $\chi_{\sigma}$. However, compared to the sigma meson, the delta meson susceptibility has a smaller overall magnitude and is less affected by the rotational effect. Since the meson susceptibilities are roughly inversely proportional to the square of the related meson masses (see, i.e., the linear sigma model result [47]), the overall magnitude of the delta meson susceptibility, associated with a heavier scalar meson, is suppressed, and the effect of the angular velocity diminishes for the heavier scalar meson.
The pseudoscalar meson susceptibilities also show comparable behavior between isotriplet and isosinglet pseudoscalar mesons. In contrast to the scalar meson susceptibilities, the overall magnitude of the pion susceptibility is suppressed with increasing the temperature, and decreases with increasing $r$ in the presence of rotation. This behavior aligns with that of the dynamical quark mass. Indeed, the pion susceptibility is directly connected to the quark condensate via the WT identity associated with the $SU(2)_{A}$ symmetry, as shown in Eq. (30). For the isosinglet channel, the eta meson susceptibility exhibits a trend similar to that of the isotriplet channel $\chi_{\pi}$, but with an overall smaller magnitude and a reduced sensitivity to the rotational effect. This is again attributed to the larger mass of the eta meson compared to the pion mass.

### v2

@@ 3.1 Meson susceptibilities
$G_{\sigma}=G_{\pi},\quad G_{\eta}=G_{\delta}=g_{s}-g_{d}.$ (3.26)
The one-loop results are given by $\chi_{\eta}^{({\rm 1-loop})}=\chi_{\pi}^{({\rm 1-loop})}$ and $\chi_{\sigma}^{({\rm 1-loop})}=\chi_{\delta}^{({\rm 1-loop})}$ with
$\begin{split}\chi_{\sigma}^{({\rm 1-loop})}\left(r;M(r)\right)&=N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\\ &\quad\times\left[\frac{p_{l,k}^{2}+p_{z}^{2}}{E^{3}}\Bigl\{f(p_{z};\Lambda)-2n_{F}(E)\Bigr\}-\frac{2M^{2}}{TE^{2}}n_{F}^{\prime}(E)\right],\end{split}$ (3.27)
where $n_{F}^{\prime}(E)=dn_{F}/dE$. As with the one-loop result for the pion susceptibility, the resumed meson susceptibilities $\chi_{\alpha}$ also become independent of rotation $\Omega$ at zero temperature. The rotational effect at finite temperature will be specifically analyzed through numerical calculations in the next section.
3.2 Other susceptibility functions
@@ 3.2.1 Topological susceptibility
${\cal L}_{\theta}=\theta Q,\quad Q=\frac{g^{2}}{64\pi^{2}}\epsilon^{\mu\nu\rho\sigma}G_{\mu\nu}^{a}G_{\rho\sigma}^{a},$ (3.28)
@@
$\int d^{4}x\chi_{\rm top}(x)=\frac{\partial^{2}\Gamma_{\rm QCD}}{\partial\theta^{2}}\Biggl|_{\theta=0}.$ (3.29)
By further evaluating the right-hand side in Eq. (3.29) with the $\theta$-term, the topological susceptibility is expressed by the two-point correlation function of the topological operator,
$\chi_{\rm top}=\int d^{4}x^{\prime}\langle T_{\tau}Q(x)Q(x^{\prime})\rangle.$ (3.30)
In the NJL model without rotation, this is evaluated as a constant form GomezNicola:2016ssy; GomezNicola:2017bhm; GomezNicola:2019myi; Kawaguchi:2020qvg; Cui:2021bqf; Cui:2022vsr: $\chi_{\rm top}=\frac{m_{0}}{4}\langle\bar{q}q\rangle+\frac{m_{0}^{2}}{4}\chi_{\eta}$. Under rotation, it becomes the radial-coordinate dependent form, as follows:
$\chi_{\rm top}\left(r;M(r)\right)=\frac{m_{0}^{2}}{4}\left[-\chi_{\pi}\left(r;M(r)\right)+\chi_{\eta}\left(r;M(r)\right)\right],$ (3.31)
where we have used the WT identity (3.7).
We should here emphasize that the topological susceptibility is a well-defined quantity even within low-energy effective theories and models which have no explicit gluon degrees of freedom. This is because the full QCD involving the $\theta$ term is mathematically equivalent to that without the $\theta$ term but with the quark mass $m\,e^{i\gamma_{5}\theta/N_{f}}$, which was pointed out in Ref. Peccei:1977hh. This equivalence is the basis in constructing the Di-Vecchia–Veneziano model DiVecchia:1980ve, which is an effective model with finite $\theta$ and provides a dynamical mechanism underlying in the Veneziano-Witten formula Witten:1979vv; Veneziano:1979ec. In the same manner, we can introduce $\theta$ in other works (see Ref. Leutwyler:1992yt for example), including the NJL model in Refs. Boer:2008ct. These effective models lead to the topological susceptibility consistent with that derived from the full QCD with the mass term $\bar{q}(m\,e^{i\gamma_{5}\theta/N_{f}})q$ GomezNicola:2016ssy; GomezNicola:2017bhm; GomezNicola:2019myi; Kawaguchi:2020qvg; Cui:2021bqf; Cui:2022vsr, and indeed shows good agreements with lattice QCD Borsanyi:2016ksw; Kawaguchi:2020qvg; Cui:2021bqf #2#2#2Indeed, the model calculations of the topological susceptibility show good agreements with the lattice QCD data in flat spacetime without rotation Borsanyi:2016ksw; Kawaguchi:2020qvg; Cui:2021bqf, as follows: 2+1+1 flavor lattice QCD (at T=0T=0Borsanyi:2016ksw: $\chi_{\rm top}=0.0245(24)_{\rm stat}(03)_{\rm flow}(12)_{\rm cont}/{\rm fm}^{4}$ 3 flavor ChPT Borsanyi:2016ksw (see also Ref. Leutwyler:1992yt): $\chi_{\text{top}}=0.0224/{\rm fm}^{4}$ 3 flavor LSM Kawaguchi:2020qvg: $\chi_{\rm top}=0.0263/{\rm fm}^{4}$ 3 flavor NJL Cui:2021bqf: $\chi_{\rm top}=0.025/{\rm fm}^{4}$, where LSM stands for the linear sigma model.. Since the $\mathrm{U(1)}_{\mathrm{A}}$ transformation is intact even under rotation, the topological susceptibility (3.31) takes a form similar to that in the nonrotating case, except for the $r$- and $\Omega$-dependence.
3.2.2 Baryon number susceptibility
@@
$\begin{split}\int d^{4}x\chi_{B}(x)&=\frac{\partial^{2}\Gamma}{\partial\mu^{2}}\Biggl|_{\mu=0}=\int d^{4}x\int d^{4}x^{\prime}\langle T_{\tau}j_{B}(x)\,j_{B}(x^{\prime})\rangle.\end{split}$ (3.32)
@@
$\chi_{B}\left(r;M(r)\right)=N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\left\{-2n_{F}^{\prime}(E)\right\}.$ (3.33)
@@ 3.2.3 Moment of inertia
$\begin{split}\int d^{4}x\chi_{\Omega}(x)&=\frac{\partial^{2}\Gamma}{\partial\Omega^{2}}\Biggl|_{\Omega=0}=\lim_{\Omega\to 0}\left[\int d^{4}x\int d^{4}x^{\prime}\langle T_{\tau}j_{\Omega}(x)\,j_{\Omega}(x^{\prime})\rangle\right].\end{split}$ (3.34)
@@
$\chi_{\Omega}\left(r;M(r)\right)=N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\left\{-2n_{F}^{\prime}(E)\right\}\left(l+\frac{1}{2}\right)^{2}.$ (3.35)
@@
4 Numerical evaluation of susceptibilities
@@
$\frac{\delta\,\Gamma\left(M\right)}{\delta M}=0.$ (4.1)
@@
$\begin{split}\Gamma\left(M\right)T&=-\int d^{3}x\frac{(M-m_{0})^{2}}{2(g_{s}+g_{D})}+N_{c}N_{f}\int d^{3}x\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\\ &\quad\times\left[Ef(p_{z};\Lambda)+2T\ln\left(1+e^{-\beta\left\{E-\Omega\left(l+\frac{1}{2}\right)\right\}}\right)\right].\end{split}$ (4.2)
@@
$M=m_{0}+M(g_{s}+g_{d})N_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\frac{f^{(l,k)}(p_{z};\Lambda)-2n_{F}(E)}{E}.$ (4.3)
@@
Using the dynamical quark mass in Eq. (4.3), one can verify the validity of the WT identity in Eq. (3.7) within the NJL model under rotation. From the dynamical quark mass, the quark condensate can be expressed as
$\langle\bar{q}q\rangle=-MN_{c}N_{f}\int\frac{dp_{z}}{2\pi}\sum_{l,k}\frac{J_{l}(p_{l,k}r)^{2}+J_{l+1}(p_{l,k}r)^{2}}{\pi R^{2}N_{l,k}^{2}}\frac{f(p_{z};\Lambda)-2n_{F}(E)}{E}.$ (4.4)
Substituting this into the pion susceptibility in Eq. (3.23), one finds that the WT identity is satisfied:
$\begin{split}\chi_{\pi}&=\frac{-\langle\bar{q}q\rangle}{M+G_{\pi}\langle\bar{q}q\rangle}=\frac{-\langle\bar{q}q\rangle}{m_{0}}.\end{split}$ (4.5)
When formulating the quark propagator within the local density approximation, we have analytically neglected the $O(\partial_{r}M/M^{2})$ term in Eq. (3.12). Therefore, our numerical analysis is valid provided that the $r$-dependent quark mass satisfies the following condition:
$\frac{|\partial_{r}M|}{M^{2}}\ll 1.$ (4.6)
@@
As we mentioned below Eq. (3.21), at $T=0$ the dynamical quark mass is unaffected by finite $\Omega$ effects, though it still exhibits $r$-dependence. To numerically examine the $r$-dependence of the dynamical quark mass in the cylindrical frame at fixed temperatures, we present panel (a) of Fig. 1. At $T=0$, the dynamical quark mass remains constant up to approximately r=0.85Rr=0.85\,R but begins to decrease sharply near the boundary to satisfy the boundary condition. Moving on to finite temperature systems, the dynamical quark mass is affected by the rotational effect even for $r<0.85R$, and it decreases with increasing $r$.
To confirm the validity of the local density approximation, we plot $|\partial_{r}M|/M^{2}$ in panel (b). This result shows that $|\partial_{r}M|/M^{2}\ll 1$ is satisfied for approximately r<0.9Rr<0.9\,R even at high temperatures across the pseudocritical temperature with the finite $\Omega$. This implies that the local density approximation is valid up to r=0.9Rr=0.9\,R even when the rotational effect is included at the high temperatures. However, in the region near the boundary for 0.9R<r<R0.9\,R<r<R, $|\partial_{r}M|/M^{2}$ increases sharply, and the validity of the local density approximation becomes unreliable. Based on this evaluation, our analysis in the following subsections will focus on the region r<0.9Rr<0.9\,R.
@@
4.1 Meson and topological susceptibilities
In Fig. 3, we numerically evaluate the meson susceptibilities under rotation. Panel (a) illustrates its $r$-dependence at fixed temperature, both with and without the rotational effect, clearly showing the difference in the $r$-dependence between the scalar meson susceptibilities $(\chi_{\sigma,\delta})$ and pseudoscalar meson susceptibilities $(\chi_{\pi,\eta})$. This difference comes from the analytical one-loop results evaluated in Eq. (3.27). Looking at the details of each meson susceptibilities, the sigma meson susceptibility, which corresponds to the isosinglet scalar channel, remains approximately constant in the relevant region (r<0.9Rr<0.9\,R) at zero temperature. At finite temperature, its magnitude becomes larger, but it still exhibits constant behavior. When including the rotational effect, the sigma meson susceptibility increases with $r$ and then decreases near the boundary in order to satisfy the boundary condition. Similarly, the delta meson susceptibility, which corresponds to the isotriplet scalar channel, shows a qualitatively similar behavior to that of the isosinglet channel $\chi_{\sigma}$. However, compared to the sigma meson, the delta meson susceptibility has a smaller overall magnitude and is less affected by the rotational effect. Since the meson susceptibilities are roughly inversely proportional to the square of the related meson masses (see, i.e., the linear sigma model result Kawaguchi:2023olk), the overall magnitude of the delta meson susceptibility, associated with a heavier scalar meson, is suppressed, and the effect of the angular velocity diminishes for the heavier scalar meson.
The pseudoscalar meson susceptibilities also show comparable behavior between isotriplet and isosinglet pseudoscalar mesons. In contrast to the scalar meson susceptibilities, the overall magnitude of the pion susceptibility is suppressed with increasing the temperature, and decreases with increasing $r$ in the presence of rotation. This behavior aligns with that of the dynamical quark mass. Indeed, the pion susceptibility is directly connected to the quark condensate via the WT identity associated with the $SU(2)_{A}$ symmetry, as shown in Eq. (3.7). For the isosinglet channel, the eta meson susceptibility exhibits a trend similar to that of the isotriplet channel $\chi_{\pi}$, but with an overall smaller magnitude and a reduced sensitivity to the rotational effect. This is again attributed to the larger mass of the eta meson compared to the pion mass.